OpenStudy (anonymous):

Add six and one-sixth plus two and two-thirds. nine and five-sixths eight and two-thirds nine and two-thirds eight and five-sixths

OpenStudy (♪chibiterasu):

So we have: \[6\frac{ 1 }{ 6 } + 2\frac{ 2 }{ 3 } = ?\]

OpenStudy (♪chibiterasu):

Well we just need to make the denominator of the second fraction 6, and we know that 2 • 3 is 6. Therefore, we must have to multiply both the top and bottom by 2 to make it be able to add with 1/6: 2/2 • 2/3 = 4/6 So now we have: \[6\frac{ 1 }{ 6 } + 2\frac{ 4 }{ 6 } = ?\] Now just add! 6 + 2 = 8 1 + 4 = 5 Using this information we can conclude that: \[6\frac{ 1 }{ 6 } + 2\frac{ 4 }{ 6 } = 8\frac{ 5 }{6}\] It can't be simplified any further without having the numerator be a decimal, so this is our final answer.

OpenStudy (anonymous):

THC Chibi

OpenStudy (♪chibiterasu):

No problem.